OpenStudy (anonymous):

is the equation true, false or open? 4y+8=6y+3

OpenStudy (anonymous):

not understanding your question..

OpenStudy (anonymous):

what is y???

OpenStudy (anonymous):

y=5/2

OpenStudy (anonymous):

4(2.5)+8=6(2.5)+3 they both equal 18 true

OpenStudy (anonymous):

true

OpenStudy (anonymous):

do you have the answer for this test ?